Find the inverse (unilateral) Laplace transforms of the following functions:

(a) \frac{2s+5}{s^{2} +5s+6}

(b) \frac{3s +5 }{s^{2} +4s+13}

(c) \frac{(s+1)^{2}}{s^{2}-s-6}

(d)\frac{5}{s^{2}(s+2)}

(e) \frac{2s+1}{(s +1)(s^{2} +2s +2)}

(f) \frac{s+2}{s(s +1)^{2}}

(g) \frac{1}{(s +1)(s +2)^{4}}

(h) \frac{s +1}{s(s +2)^{2}(s^{2} +4s+5)}

(i) \frac{s^{2}}{(s +1)^{2}(s^{2} +2s+5)}
Answer #1

(a) Given 25+5 $455+6 25+ 5 (8+2) (5+3). ... Invense Laplace transforms of 25+5 le. Ilh 25+5 ? 3275346 1 st 55+6s ( 5+55+6 (- Given 35+5 8745+13 = 3(3+2)-1sem to = 324 Cos (34) - et sin (3+). Given (5+1) 7 (5+1) 2. e) (s²_8-6) .) -575+6 = 1- 347) (5+5+6) (35+7) (3-3) (3+2) = 1+ NS- > otest2 (88) at St +16 et 2 e24 where s(t) is Dirac delta function ed Given 5 I 5(512) Let tes = 4+ siz this is the form of partial fraction decomposition. Then 5 = 5(3+2) A+ (5+2) Bt se s(s+2)= = } 동동은라